) John approached you with a business proposition to invest in a project. The Annual Cash Flows (CFt) are uncertain. However, the following probability distribution provides you with expectations:

Probable

Cash Flow (CF)

Probability P(CF)

$80,000

.25

$120,000

.65

$150,000

.10

  1. a.    What is the expected Annual Cash Flow from this project?
  2. b.   If the cash flows are expected for three years and the current rate of return on investments is 9.5%, what is the expected Present Value of the projected cash flows?
  3. c.    Would you agree to invest $275,000 in this project? Explain!!
Answer #1

Answer

  1. Expected annual cash flow = $113,000

Probable

Cash Flow (CF)

Probability P(CF)

Expected cash flow

(I)

(II)

(1) * (II)

$ 80,000

.25

$20,000

$ 120,000

.65

$78,000

$ 150,000

.10

$15,000

Total Expected Annual Cash Flow

$113,000

  1. If the cash flows are expected for three years and the current rate of return on investments is 9.5%, what is the expected Present Value of the projected cash flows?

Answer = $283,506.47

Year

Cash Flow

Present Value Factor @ 9.5%

Present Value of cash flow

(I)

(II)

(III)

(II) * (III)

1

$113,000

.9132

$103,191.60

2

$113,000

.8340

$94,242.00

3

$113,000

.7616

$86,060.80

Present Value of the projected Cash flows

$283,494.40

  1. Would you agree to invest $275,000 in this project? Explain!!

Answer – Yes , it is recommended to invest in this project

Investment decision can be taken with the help of NPV (i.e. Net present Value)

NPV can be found out using the following formula

NPV = Present Value of future cash flows – Initial investment

Present Value of future cash flows = $283,494.40

Initial investment = $275,000

Therefor NPV = $283,494.40 - $275,000

                        = $8,494.40

Since NPV is positive that means project will be profitable to the company and investment in the project is recommended.

Note: Calculation of Present value factor

Discount Factor = 1/ (1+R) N

R = Discount Rate (i.e. = 9.5%)

N = No of years

E.g. for year 2 Discount Factor = 1/ (1.095)2

                                                                = 1/ (1.095) (1.095)

                                                = 0.8340